OpenStudy (anonymous):

find the slope and the y intercept of the line 9x=10y+30

myininaya (myininaya):

write in y=mx+b form

myininaya (myininaya):

m will be slope b will be y-intercept

OpenStudy (anonymous):

y=-9/10 x+10/3 so, your slope is=-9/10

OpenStudy (anonymous):

y intercept - put x=0

OpenStudy (anonymous):

Solve for y by isolating it by subtracting 10y from both sides and subtracting 9x from both sides. You get -10y=-9x+30. Divide all terms by -10. Get y=(-9/10)+3. Yintercept is 3 and slope is (-9/10)